# Search-based DNN Testing and Retraining with GAN-enhanced Simulations

### v1

@@ IV-C2 Results
Considering feature-based distance, DESIGNATE is the best approach; indeed, it leads to the highest median () and average () values. This result is likely due to DESIGNATE integrating feature-based distance in . Further, DESIGNATE performs significantly better than other DESIGNATE variants () and has a significantly higher probability of performing better than other DESIGNATE variants (medium effect size, based on ). Expectedly, the lack of a diversity objective in DESIGNATEsingle leads to images that are less diverse than those of DESIGNATE. The lower diversity of DESIGNATEpix compared to DESIGNATE confirms that a fitness based on pixel distance is not effective to obtain feature-based diversity. Last, the lower median obtained by DESIGNATENoGAN () is likely due to the lower fidelity of simulator images compared to the ones generated by the GAN (e.g., simulator images lack details, leading to images that are more similar to each other).
Also, DESIGNATE performs significantly better, with high effect size, than TACTIC and Random.
@@
TACTIC has a feature-based distance that is similar across pairs of images, between and (avg. is ). Since in our experiments TACTIC works by modifying a set of 220 randomly selected images, and since these images present a feature-based distance of at least , it seems that the degree of diversity introduced by TACTIC through changes in weather conditions is limited. DESIGNATE, instead is more effective in creating diverse images as 75% (1st quartile) of the pairs of images generated by DESIGNATE have a distance higher than , which is higher than the best results achieved by TACTIC. Further 25% of DESIGNATE images yield a distance higher than .
@@
The random baseline leads to images with low feature-based distance, with a maximum of ; it shows that the mere sampling of the input space does not lead to identifying diverse images and that search guidance is needed.
@@
Considering pixel-based distance, all the DESIGNATE variants have the same average distance but DESIGNATE still performs better than the others, with a small to high effect size. Such results confirm our observations above for DESIGNATEsingle and DESIGNATENoGAN. Further, the comparison with DESIGNATEpix indicates that a fitness driven by pixel-based distance is less effective than a feature-based distance fitness to maximize pixel-based diversity across images. This indicates that, for the search algorithm, it is easier to generate pixel-wise diverse images when driven by a feature-based fitness. Note that augmenting feature-based distance implies changing the elements depicted in the images and, consequently, the pixel distance. This result may thus be explained by the fact that most of the changes in simulator inputs lead to pixel-based distances that are significant (i.e., above ), whereas only pixel changes that have a strong impact on the elements in the images can result into a significant feature-based distance. Consequently, feature-based distance simplifies the task of the search algorithm to explore the input space towards the search objectives. Also, DESIGNATE performs significantly better than TACTIC and Random, with small and high effect size, respectively. As per feature-based distance, Random presents lower diversity than DESIGNATE (ranges are [0.3,0.76] for Random vs. [0.14,1.00] for DESIGNATE), thus confirming that random sampling is not sufficient to identify the unsafe parts of the input space. TACTIC’s third quartile shows that only 25% of TACTIC’s image pairs have a pixel-based distance higher than 0.68 (i.e., 68% of the pixels differ) while 95% (5th percentile) of the image pairs generated by DESIGNATE present a distance higher than 0.94 (almost all the pixels differ).
@@
To summarize, DESIGNATEsingle fares better than DESIGNATE and DESIGNATEpix in terms of accuracy. However, DESIGNATE yields more diversity (both in terms of features and pixels) than DESIGNATEsingle and DESIGNATEpix. In other words, DESIGNATE and DESIGNATEsingle represent two different trade-offs and which one to adopt depends on whether accuracy or diversity is prioritized. Since DESIGNATEsingle focuses solely on minimizing DNN accuracy, without explicitly considering diversity, it is highly effective at generating images that cause DNN failures, but at the cost of producing less diverse inputs. We should note, however, that DESIGNATE is nevertheless able to detect a large number of cases where the DNN predicts poorly, as demonstrated by the zero accuracy cases representing more that 25% of generated inputs (first quartile of the distribution in Table II). Last, there is no particular reason to rely on DESIGNATEpix as it provides no advantages in terms of diversity or accuracy.
@@ IV-D RQ2: Retraining
IV-D1 Design and measurements.
We use the images generated by the approaches considered for RQ1 to retrain our case-study subject DNN. Further, to account for non-determinism, we repeat the process 10 times. For each approach and each retraining iteration, we augment the Cityscapes training set (2975 images) with the images generated in one run of the approach. We then retrain the DNN using the augmented dataset, maintaining the same configuration settings (100 epochs) as those used by the DeeplabV3 developers [10]. To avoid introducing bias due to a different number of images in the augmented datasets, for each approach, we randomly select up to 900 images from the set of images generated in RQ1. In addition to the approaches presented in RQ1 (TACTIC and random baseline), we compare DESIGNATE to the training of the DeeplabV3 model for 100 additional epochs using the whole original training set, to determine the degree of improvement achievable with a longer training time, without generating additional training images.
@@
IV-D2 Results.
Table VI provides descriptive statistics for the metric collected when testing the retrained models. Table VII provides p-values and values resulting from the comparisons of across retrained models.
The model retrained with the output of DESIGNATE has the highest median performance at , followed by DESIGNATEsingle (), DESIGNATENoGAN (), and DESIGNATEpix (). The difference between DESIGNATE and its variants is statistically () and practically () significant, which indicates that integrating a diversity objective and feature distance as fitness functions leads to better results. Further, it shows that relying on simulator images only (i.e., DESIGNATENoGAN) leads to worse results compared to retraining with GAN-generated images, thus justifying our choice.
Also, models retrained with the outputs of DESIGNATE outperform the other models. They not only improve () over the original pre-trained DeeplabV3 model (median ), but fare much better than the DeeplabV3 retrained with the cityscapes training set (, ), a set of random realistic images (, ), or TACTIC output (, ). Such results highlight that relying on changes in weather conditions (i.e., TACTIC) limits retraining and does not yield better results than what is achievable with additional training epochs. Table VII confirms that the differences between DESIGNATE and competing approaches are practically significant (i.e., ) in addition to be statistically significant (); TACTIC, instead, leads to results that do not significantly differ from the ones obtained with the original DeeplabV3 model, with a model retrained with additional epochs, or with a set of random realistic images ().
Concluding, our results show that DESIGNATE leads to the best retraining results, thus demonstrating that (1) simulator images are not as effective as realistic images to retrain DNNs and that (2) relying on GANs to generate images that simply change weather conditions is not as effective as combining GAN with a simulator to generate realistic images for DNN retraining.
@@
Pre-trained DeeplabV3 | Retrained DeeplabV3 | DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | |||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | VDA | p-value | ||||||||
Pre-trained DeeplabV3 | 0.47 | 0.06 | 0.44 | 0.00E+00 | 0.35 | 0.00E+00 | 0.46 | 0.00E+00 | 0.44 | 0.00E+00 | 0.46 | 0.00E+00 | 0.47 | 2.00E-02 | ||
Retrained DeeplabV3 | 0.53 | 6.00E-02 | 0.47 | 0.00E+00 | 0.37 | 0.00E+00 | 0.49 | 5.00E-02 | 0.46 | 0.00E+00 | 0.49 | 6,00E-02 | 0.49 | 5.00E-01 | ||
DESIGNATEsingle | 0.56 | 0.00E+00 | 0.53 | 0.00 | 0.40 | 0.00E+00 | 0.52 | 0.00E+00 | 0.49 | 0.00E+00 | 0.52 | 0,00E+00 | 0.53 | 0.00E+00 | ||
DESIGNATE | 0.65 | 0.00E+00 | 0.63 | 0.00 | 0.60 | 0.00E+00 | 0.61 | 0.00E+00 | 0.59 | 0.00E+00 | 0.61 | 0,00E+00 | 0.62 | 0.00E+00 | ||
DESIGNATEpix | 0.54 | 0.00E+00 | 0.51 | 0.05 | 0.48 | 0.00E+00 | 0.39 | 0.00E+00 | 0.48 | 0.00E+00 | 0.50 | 8,40E-01 | 0.51 | 1.40E-01 | ||
DESIGNATENoGAN | 0.56 | 0.00E+00 | 0.54 | 0.00E+00 | 0.51 | 1.00E-01 | 0.41 | 0.00E+00 | 0.52 | 0.00E+00 | 0.53 | 0.00E+00 | 0.53 | 0.00E+00 | ||
Random | 0.54 | 0.00E+00 | 0.51 | 6.00E-02 | 0.48 | 0.00E+00 | 0.39 | 0,00E+00 | 0,5 | 8,40E-01 | 0.47 | 0.00E+00 | 0.51 | 1.90E-01 | ||
TACTIC | 0.53 | 2.00E-02 | 0.51 | 0.50 | 0.47 | 0.00E+00 | 0.38 | 0.00E+00 | 0.49 | 1.40E-01 | 0.47 | 0.00E+00 | 0.49 | 1,90E-01 |

### v2

@@ IV-C2 Results
Considering feature-based distance, DESIGNATE stands out with the highest median (14.97) and average (15.25) values. This result is likely due to DESIGNATE integrating feature-based distance in . Further, DESIGNATE performs significantly better than other DESIGNATE variants () and has a significantly higher probability of performing better than other DESIGNATE variants (medium effect size, based on ). Expectedly, the lack of a diversity objective in DESIGNATEsingle leads to images that are less diverse than those of DESIGNATE. The lower diversity of DESIGNATEpix compared to DESIGNATE confirms that a fitness based on pixel distance is not effective to obtain feature-based diversity. Last, the lower median obtained by DESIGNATENoGAN () is likely due to the lower fidelity of simulator images compared to the ones generated by the GAN (e.g., simulator images lack details, leading to images that are more similar to each other).
@@
DESIGNATE performs significantly better, with high effect size, than TACTIC and Random. TACTIC has a feature-based distance that is similar across pairs of images, between and (avg. is ). Since in our experiments TACTIC works by modifying a set of 220 randomly selected images, and since these images present a feature-based distance of at least , it seems that the degree of diversity introduced by TACTIC through changes in weather conditions is limited. DESIGNATE, instead is more effective in creating diverse images as 75% (1st quartile) of the pairs of images generated by DESIGNATE have a distance higher than , which is higher than the best results achieved by TACTIC. Further 25% of DESIGNATE images yield a distance higher than .
@@
The random baseline leads to images with low feature-based distance, with a median of and a maximum of ; it shows that the mere sampling of the input space does not lead to identifying diverse images and that search guidance is needed.
DESIGNATE significantly outperforms DeepJanus and DeepJanusGAN. Their lower diversity suggests that their search algorithm, which focuses on identifying the boundaries of correctly processed inputs, is unexpectedly less effective than NSGA-II in exploring the search landscape. DeepJanusGAN achieves higher diversity values () than DeepJanus thus confirming that GANs, by avoiding unrealistic cases (e.g., cars with few visible pixels, as discussed above), prevents the search from getting stuck in unrealistic, similar images.
@@
Considering pixel-based distance, all the DESIGNATE variants have the same average distance but DESIGNATE still performs better than the others, with a small to high effect size. Such results confirm our observations above for DESIGNATEsingle and DESIGNATENoGAN. Further, the comparison with DESIGNATEpix indicates that a fitness driven by pixel-based distance is less effective than a feature-based distance fitness to maximize pixel-based diversity across images. This indicates that, for the search algorithm, it is easier to generate pixel-wise diverse images when driven by a feature-based fitness. Note that augmenting feature-based distance implies changing the elements depicted in the images and, consequently, the pixel distance. This result may thus be explained by the fact that most of the changes in simulator inputs lead to pixel-based distances that are significant (i.e., above ), whereas only pixel changes that have a strong impact on the elements in the images can result into a significant feature-based distance. Consequently, feature-based distance simplifies the task of the search algorithm to explore the input space towards the search objectives. Also, DESIGNATE performs significantly better than TACTIC and Random, with small and high effect size, respectively. As per feature-based distance, Random presents lower diversity than DESIGNATE (ranges are [0.3,0.76] for Random vs. [0.14,1.00] for DESIGNATE), thus confirming that random sampling is not sufficient to identify the unsafe parts of the input space. TACTIC’s third quartile shows that only 25% of TACTIC’s image pairs have a pixel-based distance higher than 0.68 (i.e., 68% of the pixels differ) while 95% (5th percentile) of the image pairs generated by DESIGNATE present a distance higher than 0.94 (almost all the pixels differ). ††margin: 2.1 Last, the pixel-based distance of DESIGNATE images is significantly larger than that of DeepJanus and DeepJanusGAN images, with medium to large effect sizes, thus confirming the differences discussed above.
In the Martian environment, DESIGNATE achieved a median diversity of 14.75 for feature-based distance that is significantly higher than DESIGNATEsingle (10.59), DESIGNATEpix (11.92), the random baseline (11.36), DeepJanus (12.81), and DeepJanusGAN (9.52). However, DESIGNATENoGAN outperformed DESIGNATE in terms of feature-based diversity, which we attribute to DESIGNATENoGAN relying on the lower-fidelity outputs of the simulator; indeed, such simulator images are likely to cause erroneous outputs and therefore DESIGNATENoGAN is likely to generate images that are spread throughout the whole input space, thus leading to larger diversity. In all likelihood, this phenomenon is not observed with DeepJanus because, by generating images on the boundaries of the input space, it likely leads to more similar images (e.g., images containing roads where cars can be positioned far away).
For pixel-based diversity, DESIGNATE outperformed all the other approaches with a median pixel distance of 1.00. Instead, with feature-based distance, DESIGNATENoGAN outperformed DESIGNATE, such a difference is likely due to the nature of the two distance metrics; indeed, the pixel distance may capture dissimilarities that are practically irrelevant. For example, two images may have slightly different luminosity, thus leading to high pixel-based distance, but contain the same objects thus yielding low feature-based distance.
@@
To summarize, DESIGNATEsingle fares better than DESIGNATE and DESIGNATEpix in terms of accuracy. However, DESIGNATE yields more diversity (both in terms of features and pixels) than DESIGNATEsingle and DESIGNATEpix. In other words, DESIGNATE and DESIGNATEsingle represent two different trade-offs and which one to adopt depends on whether accuracy or diversity is prioritized. Since DESIGNATEsingle focuses solely on minimizing DNN accuracy, without explicitly considering diversity, it is highly effective at generating images that cause DNN failures, but at the cost of producing less diverse inputs. We should note, however, that DESIGNATE is nevertheless able to detect a large number of cases where the DNN predicts poorly, as demonstrated by the zero accuracy cases representing more that 25% of generated inputs (first quartile of the distribution in Table II). Last, there is no particular reason to rely on DESIGNATEpix as it provides no advantages in terms of diversity or accuracy. With respect to other alternative techniques, as presented above, they tend to either not be effective at triggering failures or generate many unrealistic and irrelevant test inputs.
@@ IV-D RQ2: Retraining
IV-D1 Design and measurements
We use the images generated by the approaches considered for RQ1 to retrain our case-study subject DNN. Further, to account for non-determinism, we repeat the process 10 times. For each approach and each retraining iteration, we use the images generated in one run of the approach to augment the Cityscapes training set (2975 images), ††margin: R3.6 for the urban environment, and the AI4MARS training set (16064 images), for the Martian environment. We then retrain the DNN using the augmented dataset, maintaining the same configuration settings (100 epochs) as those used by the DeeplabV3 developers [11]. To avoid introducing bias due to a different number of images in the augmented datasets, for each approach, we randomly select up to 900 images from the set of images generated in RQ1. In addition to the approaches presented in RQ1 (TACTIC, random baseline, DeepJanus, and DeepJanusGAN), we compare DESIGNATE to the training of the DeeplabV3 model for 100 additional epochs using the whole original training set, to determine the degree of improvement achievable with a longer training time, without generating additional training images.
@@ IV-D1 Design and measurements
For the Martian environment, we measure the obtained by all the retrained DNNs on the AI4MARS test set (966 images).
IV-D2 Results
††margin: R3.6Tables X and XII provide descriptive statistics for and from testing the retrained models in the urban and Martian environments, respectively. These tables summarize the performance improvements achieved through retraining with outputs generated by different approaches. Tables XI and XIII present the p-values and values from statistical comparisons of and across retrained models, capturing the significance and effect sizes of the observed differences.
For the urban environment, the DNN retrained with the output of DESIGNATE has the highest median performance at , followed by DeepJanus (), DESIGNATEsingle (), DeepJanusGAN (), DESIGNATENoGAN (), and DESIGNATEpix (). DESIGNATE not only improves () over the original pre-trained DeeplabV3 model (median ), but fare much better than the DeeplabV3 retrained with the cityscapes training set (, ), a set of random realistic images (, ), TACTIC (, ), DeepJanus (, ) and DeepJanusGAN (, ) outputs.
The difference between DESIGNATE and its variants is statistically () and practically () significant, which indicates that integrating a diversity objective and feature distance as fitness functions leads to better results.
Also the difference between DESIGNATE and the two DeepJanus variants is statistically and practically significant. The fact that DESIGNATE outperforms DeepJanusGAN indicates that focusing on corner cases (i.e., the frontier of behaviours found by DeepJanusGAN) is not effective to retrain a DNN. Further, the likelihood of DeepJanus and DeepJanusGAN performing better than random is very low ( and ).
In general, we can conclude that relying on simulator images only (i.e., DESIGNATENoGAN and DeepJanus) leads to worse results compared to retraining with GAN-generated images, thus justifying our choice.
Further, models retrained relying on search to drive the generation of scenarios (DESIGNATE and DeepJanus variants) outperform the other approaches. Such results suggest that relying on changes in weather conditions (i.e., TACTIC) limits retraining and does not yield better results than what is achievable with additional training epochs. Further, TACTIC leads to results that do not significantly differ from the ones obtained with the original DeeplabV3 model, with a model retrained with additional epochs, or with a set of random realistic images ().
For the Martian environment, DESIGNATE again demonstrated superior performance, achieving the highest median (0.53, +0.09 over the original model). These results highlight the effectiveness of DESIGNATE in generating both failure-inducing and diverse inputs, which, when used for retraining, improves DNN robustness in challenging environments. The difference between DESIGNATE and its variants is statistically () and practically () significant, thus confirming the effectiveness of integrating a diversity objective and feature distance as fitness functions.
Like in the urban environment, DESIGNATE () outperforms DeepJanus () and DeepJanusGAN (), thus confirming to be a better choice for retraining. Also, DESIGNATE outperforms a random approach (), with a practically significant difference (), while DeepJanus and DeepJanusGAN perform similar to random ().
Concluding, our results show that DESIGNATE leads to the best retraining results, thus demonstrating that (1) simulator images are not as effective as realistic images to retrain DNNs, ††margin: R3.6 (2) relying on images belonging to the frontier of DNN behaviours is not the best approach for DNN retraining, and (3) relying on GANs to generate images that simply change weather conditions is not as effective as combining GAN with a simulator to generate realistic images for DNN retraining.
@@ IV-D2 Results
DeepJanus | 0.00 (+0.00) | 0.98 (+0.01) | 0.87 (+0.05) | 0.00 (+0.00) | 0.75 (+0.13) | 0.93 (+0.02) | 0.77 (+0.06) |
DeepJanusGAN | 0.00 (+0.00) | 0.97 (+0.00) | 0.86 (+0.04) | 0.00 (+0.00) | 0.73 (+0.11) | 0.92 (+0.01) | 0.76 (+0.05) |
Note: best (i.e., highest or most diverse) results (per column) in bold.
@@
Pre-trained DeeplabV3 | Retrained DeeplabV3 | DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | DeepJanus | DeepJanusGAN | |||||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | VDA | p-value | VDA | p-value | VDA | p-value | ||||||||
Pre-trained DeeplabV3 | 0.47 | 0.06 | 0.44 | 0.00E+00 | 0.35 | 0.00E+00 | 0.46 | 0.00E+00 | 0.44 | 0.00E+00 | 0.46 | 0.00E+00 | 0.47 | 2.00E-02 | 0.42 | 0.00E+00 | 0.44 | 0.00E+00 | ||
Retrained DeeplabV3 | 0.53 | 6.00E-02 | 0.47 | 0.00E+00 | 0.37 | 0.00E+00 | 0.49 | 5.00E-02 | 0.46 | 0.00E+00 | 0.49 | 6,00E-02 | 0.49 | 5.00E-01 | 0.44 | 0.00E+00 | 0.47 | 0.00E+00 | ||
DESIGNATEsingle | 0.56 | 0.00E+00 | 0.53 | 0.00 | 0.40 | 0.00E+00 | 0.52 | 0.00E+00 | 0.49 | 0.00E+00 | 0.52 | 0,00E+00 | 0.53 | 0.00E+00 | 0.47 | 0.00E+00 | 0.50 | 7.00E-01 | ||
DESIGNATE | 0.65 | 0.00E+00 | 0.63 | 0.00 | 0.60 | 0.00E+00 | 0.61 | 0.00E+00 | 0.59 | 0.00E+00 | 0.61 | 0,00E+00 | 0.62 | 0.00E+00 | 0.57 | 0.00E+00 | 0.59 | 0.00E+00 | ||
DESIGNATEpix | 0.54 | 0.00E+00 | 0.51 | 0.05 | 0.48 | 0.00E+00 | 0.39 | 0.00E+00 | 0.48 | 0.00E+00 | 0.50 | 8,40E-01 | 0.51 | 1.40E-01 | 0.45 | 0.00E+00 | 0.48 | 0.00E+00 | ||
DESIGNATENoGAN | 0.56 | 0.00E+00 | 0.54 | 0.00E+00 | 0.51 | 1.00E-01 | 0.41 | 0.00E+00 | 0.52 | 0.00E+00 | 0.53 | 0.00E+00 | 0.53 | 0.00E+00 | 0.48 | 0.00E+00 | 0.51 | 2.70E-01 | ||
Random | 0.54 | 0.00E+00 | 0.51 | 6.00E-02 | 0.48 | 0.00E+00 | 0.39 | 0,00E+00 | 0.50 | 8,40E-01 | 0.47 | 0.00E+00 | 0.51 | 1.90E-01 | 0.45 | 0.00E+00 | 0.48 | 0.00E+00 | ||
TACTIC | 0.53 | 2.00E-02 | 0.51 | 0.50 | 0.47 | 0.00E+00 | 0.38 | 0.00E+00 | 0.49 | 1.40E-01 | 0.47 | 0.00E+00 | 0.49 | 1,90E-01 | 0.44 | 0.00E+00 | 0.47 | 0.00E+00 | ||
DeepJanus | 0.58 | 0.00E+00 | 0.56 | 0.00E+00 | 0.53 | 0.00E+00 | 0.43 | 0.00E+00 | 0.55 | 0.00E+00 | 0.52 | 0.00E+00 | 0.55 | 0.00E+00 | 0.56 | 0.00E+00 | 0.53 | 0.00E+00 | ||
DeepJanusGAN | 0.56 | 0.00E+00 | 0.53 | 0.00E+00 | 0.50 | 7.00E-01 | 0.41 | 0.00E+00 | 0.52 | 0.00E+00 | 0.49 | 2.70E-01 | 0.52 | 0.00E+00 | 0.53 | 0.00E+00 | 0.47 | 0.00E+00 |
Note: tangible (based on ) differences in bold; for each pair of approaches, the best approach is the one with and its name on the row.
None (Original DeeplabV3) | 0.00 | 0.99 | 0.44 | 0.01 | 0.26 | 0.77 | 0.49 |
AI4Mars training (Retrained DNN) | 0.01 (+0.01) | 1.00 (+0.01) | 0.47 (+0.03) | 0.17 (+0.16) | 0.33 (+0.07) | 0.63 (-0.14) | 0.49 (+0.00) |
DESIGNATEsingle | 0.00 (+0.00) | 1.00 (+0.01) | 0.50 (+0.06) | 0.19 (+0.18) | 0.34 (+0.08) | 0.67 (-0.10) | 0.52 (+0.03) |
DESIGNATE | 0.00 (+0.00) | 1.00 (+0.01) | 0.53 (+0.09) | 0.19 (+0.18) | 0.38 (+0.12) | 0.69 (-0.08) | 0.53 (+0.04) |
DESIGNATEpix | 0.00 (+0.00) | 1.00 (+0.01) | 0.50 (+0.06) | 0.15 (+0.14) | 0.36 (+0.09) | 0.68 (-0.09) | 0.52 (+0.02) |
DESIGNATENoGAN | 0.00 (+0.00) | 1.00 (+0.01) | 0.46 (+0.02) | 0.07 (+0.06) | 0.30 (+0.04) | 0.64 (-0.13) | 0.47 (-0.02) |
Random | 0.00 (+0.00) | 1.00 (+0.01) | 0.49 (+0.05) | 0.12 (+0.10) | 0.33 (+0.07) | 0.68 (-0.09) | 0.51 (+0.02) |
DeepJanus | 0.00 (+0.00) | 1.00 (+0.01) | 0.50 (+0.06) | 0.19 (+0.18) | 0.36 (+0.09) | 0.66 (-0.10) | 0.52 (+0.03) |
DeepJanusGAN | 0.00 (+0.00) | 1.00 (+0.01) | 0.50 (+0.06) | 0.20 (+0.19) | 0.36 (+0.09) | 0.67 (-0.10) | 0.52 (+0.03) |
Note: best (i.e., highest or most diverse) results (per column) in bold.
Original
Original_retrained
DESIGNATEsingle
DESIGNATE
DESIGNATEpix
DESIGNATENoGAN
Random
DeepJanus
DeepJanusGAN
p-value
p-value
p-value
p-value
p-value
p-value
p-value
p-value
Original
0.47
1.60E-01
0.47
1.00E-02
0.39
0.00E+00
0.43
0.00E+00
0.51
5.70E-01
0.48
1.60E-01
0.44
0.00E+00
0.44
0.00E+00
Retrained DNN
0.53
1.60E-01
0.50
7.40E-01
0.44
0.00E+00
0.48
2.30E-01
0.54
0.00E+00
0.52
1.80E-01
0.49
3.90E-01
0.49
3.10E-01
DESIGNATEsingle
0.53
1.00E-02
0.50
7.40E-01
0.43
0.00E+00
0.47
0.00E+00
0.54
0.00E+00
0.51
3.00E-02
0.48
0.00E+00
0.48
0.00E+00
DESIGNATE
0.61
0.00E+00
0.56
0.00E+00
0.57
0.00E+00
0.54
0.00E+00
0.61
0.00E+00
0.58
0.00E+00
0.55
0.00E+00
0.55
0.00E+00
DESIGNATEpix
0.57
0.00E+00
0.52
2.30E-01
0.53
0.00E+00
0.46
0.00E+00
0.57
0.00E+00
0.54
0.00E+00
0.51
1.00E-02
0.51
5.00E-02
DESIGNATENoGAN
0.49
5.70E-01
0.46
0.00E+00
0.46
0.00E+00
0.39
0.00E+00
0.43
0.00E+00
0.47
0.00E+00
0.44
0.00E+00
0.44
0.00E+00
Random
0.52
1.60E-01
0.48
1.80E-01
0.49
3.00E-02
0.42
0.00E+00
0.46
0.00E+00
0.53
0.00E+00
0.47
0.00E+00
0.47
0.00E+00
DeepJanus
0.56
0.00E+00
0.51
3.90E-01
0.52
0.00E+00
0.45
0.00E+00
0.49
1.00E-02
0.56
0.00E+00
0.53
0.00E+00
0.50
6.30E-01
DeepJanusGAN
0.56
0.00E+00
0.51
3.10E-01
0.52
0.00E+00
0.45
0.00E+00
0.49
5.00E-02
0.56
0.00E+00
0.53
0.00E+00
0.50
6.30E-01
Note: tangible (based on ) differences in bold; for each pair of approaches, the best approach is the one with and its name on the row.